OpenStudy (moongazer):

Determine the equation of the line in general form given the following condition. 2.) P(-1,5) m=0

OpenStudy (moongazer):

is the answer y-5=0 ?

OpenStudy (moongazer):

I'm just checking

OpenStudy (anonymous):

is it m a gradient?

OpenStudy (moongazer):

slope

OpenStudy (anonymous):

then we know all along the equation of line is y=mx+c where c is a constant, then with information P(-1,5), then 5=(0)(-1)+c c=5 the line are y=5 then. a straight line

OpenStudy (moongazer):

I know it is a straight line I just want to know if my equation is correct y-5=0 is this the general form of the line?

OpenStudy (anonymous):

yeah u can say y-5=0 or y=5, they are equally same

OpenStudy (moongazer):

thanks

OpenStudy (moongazer):

I am just confirming my answer. :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

:)